In a router for scattering experiment and alpha particle is positive charge 2 is directed towards a gold necklace with charms 79-80 alpha particle has a kinetic energy of 5.0 even went very far from the nucleus as even the Golden close to be fixed and space determine the distance of the closest approach note 8.99 * 10 is 4 and 9 Newton per hour metre square column column
1
Expert's answer
2020-03-16T13:15:51-0400

From the conservation of energy:


"K=\frac{kq_1q_2}{r}"


"5(1.6\cdot 10^{-13})=\frac{(8.99\cdot 10^{9})(2)(79)(1.6\cdot 10^{-19})^2}{r}"

"r=4.5\cdot 10^{-14}m"